The vapor pressure of benzene (C

6

H

6

) is 40.1 mmHg at 7.6°C. What is its vapor pressure at

60.6°C? The molar heat of vaporization of benzene is 31.0 kJ/mol


1
Expert's answer
2020-08-26T03:08:12-0400

Clausius-Clapeyron equation:

"ln(\frac{p_2}{p_1}) = (\frac{∆H_{vap}}{R})\times(\frac{1}{T_1} – \frac{1}{T_2})"

R — the gas constant (8.3145 J/mol K)

p1 – the vapor pressure of benzene at T1;

p2 – the vapor pressure of benzene at T2;

∆Hvap – the enthalpy of vaporization;

p1 = 40.1 mmHg

T1 = 7.6 °C + 273.15 = 280.75 K

T2 = 60.6 °C + 273.15 = 333.75 K

∆Hvap = 31000 J/mol

"ln(\frac{p_2}{40.1}) = \frac{31000}{8.3145}\times(\frac{1}{280.75} – \frac{1}{333.75})"

"ln(\frac{p_2}{40.1}) = 3728.42×(0.00356 – 0.00299)"

"ln(\frac{p_2}{40.1}) = 3728.42×0.00057"

"ln(\frac{p_2}{40.1}) = 2.12"

"\frac{p_2}{40.1} = e^{2.12} = 8.33"

"p_2 = 40.1\times8.33 = 334.03 mmHg"

Answer: 334.03 mmHg